# DeepTV: A neural network approach for total variation minimization

### v1

@@ 1 Introduction
$\operatorname{TV}(u):=\sup\left\{\int_{\Omega}u\operatorname*{div}\vec{q}\ \text{d}x\colon\vec{q}\in C_{c}^{1}(\Omega)^{d},|\vec{q}(x)|_{2}\leq 1\text{ for allmost all }x\in\Omega\right\},$ (2)
@@
We choose to replace the solution space $BV(\Omega)\cap L^{2}(\Omega)$ in (1) by a set of neural networks $\mathcal{H}$, whose activation functions are ReLUs (rectifier linear units). In the sequel we will call such neural networks ReLU neural networks (ReLU-NNs). While any other type of neural networks might be used in this approach too, we note that ReLU-NNs are theoretically quite well studied, which is useful for our theoretical analysis. The resulting optimization problem reads then
$\inf_{u\in\mathcal{H}}\alpha_{1}\|Tu-g\|_{L^{1}(\Omega)}+\alpha_{2}\|Tu-g\|_{L^{2}(\Omega)}^{2}+\lambda\operatorname{TV}(u).$ (3)
@@
As the Deep Ritz Method and PINNs, this approach can be interpreted as unsupervised learning. However, in a more classical view (3) is simply an optimization problem with solution space $\mathcal{H}$. Unfortunately, it turns out that even if (1) has a solution in $BV(\Omega)\cap L^{2}(\Omega)$ this solution is in general not in $\mathcal{H}$, see example 3.3 below. This has also been observed in [9]. The reason for this negative result is that the space $BV(\Omega)\cap L^{2}(\Omega)$ contains functions with jumps and ReLU-NNs cannot represent such functions. In fact ReLU-NNs can “only” represent continuous and piecewise affine functions [7]. Despite this negative observation, we suggest an auxiliary neural network model, which is a constrained version of (3). That is we add a constraint to (3) such that a solution in $\mathcal{H}$ is indeed guaranteed. We prove that this auxiliary model $\Gamma$-converges to the $L^{1}$-$L^{2}$-TV model, which fully justifies our approach. We are aware that $\Gamma$-convergence for a Deep Ritz Method is also studied in [21], but in a completely different setting, i.e. for a different type of optimization problem, namely for variational problems arising from partial differential equations, and in Sobolev spaces and not in the space of bounded variation. Hence our analysis completely differs from the one in [21]. However, our $\Gamma$-convergence proof is inspired by the $\Gamma$-convergence analysis in [34], where the $\Gamma$-convergence of a discrete total variation to the continuous total variation (2) is presented.
@@
To numerically compute a solution of the auxiliary model, numerical quadrature to approximate the integrals are needed. Note that since $\mathcal{H}\subset W^{1,1}(\Omega)$, we have that $\operatorname{TV}(u)=\int_{\Omega}|\nabla u|_{2}\text{d}x$ for any $u\in\mathcal{H}$. Therefore, the total variation of $u\in\mathcal{H}$ is computed using the gradient of $u$, which must be carefully approximated in a numerical setting, i.e. when applying quadrature. It seems reasonable to approximate the gradient by a finite difference approximation. The reason for this is that replacing the integral by a numerical quadrature evaluates the gradient only at a finite number of points. However, this can be problematic for piecewise constant functions, because the gradient may be zero at the specific evaluation points chosen by the quadrature, even though the overall function might have significant variation. This leads to an inaccurate computation of the total variation. A finite difference approximation of the gradient circumvents this shortcoming. We would like to emphasize that despite the described discretizations we still optimize the resulting discrete auxiliary model over the space $\mathcal{H}$. Further, we are able to show that this discrete auxiliary model again $\Gamma$-converges to (1). In particular, the $\Gamma$-convergence proof suggests a particular discretization of the gradient and hence total variation which ensures that the discrete solutions approximate the continuous solution as the discretization becomes finer. The same particular discretization of the total variation has been already used in [34] to show $\Gamma$-convergence and later in [33], but not in the context of neural networks. Our proof is actually based on the $\Gamma$-convergence proof in [34] and adjusted to our neural network setting. However, any convergence result of minimizers of the discrete auxiliary model to a solution of the continuous problem is missing in [34], while we present such a result in our setting. In [33] a convergence result of minimizers is presented, however with respect to the $L^{2}(\Omega)$ topology and under certain regularity assumption on the data $g$, which we do not need. Hence our result is more general and even stated with respect to the $L^{1}(\Omega)$ topology.
@@
The rest of the paper is organized as follows: In section 2, we review and introduce essential definitions related to neural networks and functional analysis, which are foundational for the subsequent sections. We analyze the neural network problem (3) in section 3 and in particular suggest an auxiliary model which indeed possesses a minimizer in $\mathcal{H}$. We show its $\Gamma$-convergence to the $L^{1}$-$L^{2}$-TV model and provide an error estimate. Moreover, we present a universal approximation result for $BV$-functions. In section 4 we describe the numerical realization of the proposed model. In particular, we present a numerical realization of the total variation. We connect the fully discretized neural network model to the original model (1) by utilizing $\Gamma$-convergence. Further we discuss a numerical issue and propose a solution to circumvent it. In section 5 we present several numerical experiments supporting our theoretical findings. We conclude with a short discussion in section 6.
@@ 2 Definitions and notations
As we are interested to represent functions $u:\Omega\to\mathbb{R}$ with a ReLU-NN, we consider only ReLU-NNs whose input layer consists of $d\in\{1,2\}$ neurons and whose output layer consists of 1 neuron. Thereby the number of neurons in the hidden layers might be arbitrary. We denote the set of ReLU-NNs with this fixed architecture by $\mathcal{H}$. The collection of all weights and biases of a neural network in $\mathcal{H}$ is denoted by $\theta\in\mathbb{R}^{M}$, $M\in\mathbb{N}$. That is the sum of the number of weights and biases of a neural network $u_{\theta}\in\mathcal{H}$ is $M$. For a neural network $u_{\theta}\in\mathcal{H}$ we use the index $\theta$ to emphasize that the neural network depends on the weights and biases $\theta$.
@@ 3 Problem description
$\min_{u_{\theta}\in\mathcal{H}}\alpha_{1}\|Tu_{\theta}-g\|_{L^{1}(\Omega)}+\alpha_{2}\|Tu_{\theta}-g\|_{L^{2}(\Omega)}^{2}+\lambda\operatorname{TV}(u_{\theta}).$ (4)
@@ 3.1 Existence of a minimizer
Let $\mathcal{H}$ be the set of one hidden layer neural networks, where each one hidden layer neural network is defined as
@@
Here $m\in\mathbb{N}$ is the number of neurons in the hidden layer, $w\in\mathbb{R}^{m}$ is the vector of weights from the input neuron to the hidden layer, $a\in\mathbb{R}^{m}$ is the vector of weights from the hidden layer to the output neuron, $b\in\mathbb{R}^{m}$ is a vector of biases, and $\sigma:\mathbb{R}\to\mathbb{R}$ is the ReLU activation function, i.e. $\sigma(x)=\max\{0,x\}$. If we try to approximate the BV-function
@@
on $\Omega=(-l_{\ell},l_{u})\subset\mathbb{R}$ by solving $\min_{u_{w,a,b}\in\mathcal{H}}\|u_{w,a,b}-g\|_{L^{2}(\Omega)}^{2}$, i.e. we set $T=I$, $\alpha_{1}=0=\lambda$ and $\alpha_{2}=1$ in (4), then it turns out that this problem does not have a solution. In fact, any function in $\mathcal{H}$ is continuous and since $g$ is not continuous we obtain that $\|u_{w,a,b}-g\|_{L^{2}(\Omega)}^{2}>0$ for all $u_{w,a,b}\in\mathcal{H}$. To see that this is not a contradiction to the universal approximation theorem, see theorems 3.1 and 3.2, i.e. that any $BV$-function can be approximated by a ReLU-NN arbitrarily good, let us set $m=2$, $a_{1}=1$ and $a_{2}=-1$. Further we choose $w_{1}=\frac{1}{h}=w_{2}$, $h>0$, and $b_{1}=1$, $b_{2}=0$. This yields
@@
where $c_{1},c_{2}\in\mathbb{R}^{+}_{0}$ such that $c_{1}\leq c_{2}$ and $c_{1},c_{2}$ depend on $\alpha_{1}$ and $\alpha_{2}$. The analytic computation of the exact solution for this example can be founded in appendix A. Hence, if $\alpha_{1}$ and $\alpha_{2}$ are chosen such that $c_{1}<c_{2}$, then a function in $\mathcal{H}$ cannot solve this problem, but only approximates a solution arbitrary close. From example 3.3 we learn that the failure of the existence of a solution of (4) might be due to the unboundedness of the weights $\theta$, rendering the solution space not compact. Thus we slightly modify (4) to
$\min_{u_{\theta}\in\mathcal{H}_{c}}\alpha_{1}\|Tu_{\theta}-g\|_{L^{1}(\Omega)}+\alpha_{2}\|Tu_{\theta}-g\|_{L^{2}(\Omega)}^{2}+\lambda\operatorname{TV}(u_{\theta})$ (9)
where $\mathcal{H}_{c}:=\{u_{\theta}\in\mathcal{H}\colon|\theta|_{\infty}\leq c\}$ with $c\geq 0$ typically to be chosen large such that (9) is a close approximation of (4). Note that in $\mathbb{R}^{M}$ all norms are equivalent and hence any arbitrary vector-norm could be chosen instead of $|\cdot|_{\infty}$.
@@
The optimization problem in (9) possesses a solution $u_{\theta}\in\mathcal{H}$.
@@
where $\Theta:=\{\theta\in\mathbb{R}^{M}\colon|\theta|_{\infty}\leq c\}$. Since $\Theta$ is closed and bounded, it is compact (Heine-Borel Theorem). Let us denote the functional in (10) by $J(\theta):=J_{c}(\theta)+J_{d}(\theta)$, where $J_{c}(\theta):=\alpha_{1}\|Tu_{\theta}-g\|_{L^{1}(\Omega)}+\alpha_{2}\|Tu_{\theta}-g\|_{L^{2}(\Omega)}^{2}$ and $J_{d}(\theta):=\lambda\operatorname{TV}(u_{\theta})$. We show now that $J$ is lower semi-continuous. Since the ReLU activation function $\sigma$ is continuous, functions in $\mathcal{H}$ are continuous with respect to $\theta$. Hence, for a sequence $(\theta_{n})_{n}\subset\mathbb{R}^{M}$ with limit $\theta\in\mathbb{R}^{M}$ for $n\to\infty$ we have $u_{\theta_{n}}\to u_{\theta}$ as $\theta_{n}\to\theta$. Since strong convergence implies weak convergence we also have $u_{\theta_{n}}\rightharpoonup u_{\theta}$ as $\theta_{n}\to\theta$. By the weak lower semi-continuity of the total variation we obtain $\operatorname{TV}(u_{\theta})\leq\liminf_{\theta_{n}\to\theta}\operatorname{TV}(u_{\theta_{n}})$ and hence $J_{d}$ is lower semi-continuous. Further, since $J_{c}$ is continuous, we have that $J$ is lower semi-continuous.
@@
We note that theorem 3.4 can be extended to any activation function $\sigma\in C(\mathbb{R})$, as the proof only requires that the activation functions are continuous. That is if in (9) the solution space $\mathcal{H}$ is replaced by a set of neural networks with arbitrary continuous activation functions, then (9) still possesses a solution in this respective space.
@@
$\min_{u_{\theta}\in\mathcal{H}}\alpha_{1}\|Tu_{\theta}-g\|_{L^{1}(\Omega)}+\alpha_{2}\|Tu_{\theta}-g\|_{L^{2}(\Omega)}^{2}+\lambda\operatorname{TV}(u_{\theta})+\alpha_{\theta}|\theta|_{\infty},$ (11)
@@
Note that theorem 3.4 guarantees “only” the existence of a minimizer, but does not say anything about the uniqueness. In fact, it seems elusive to show uniqueness of a solution of (9), since $u_{\theta}$ is usually non-convex with respect to $\theta$ rendering (9) non-convex. On the contrary (1) is convex with respect to $u$ and even strictly convex if $\alpha_{2}>0$ and $T$ is injective. This means that changing the solution space in (1) to $\mathcal{H}$ or $\mathcal{H}_{c}$ also changes the nature of the problem. Still (1) and (9) are closely related as we will show in the next subsection.

### v2

@@ 1 Introduction
$\operatorname{TV}(u):=\sup\left\{\int_{\Omega}u\operatorname*{div}q\ \text{d}x\colon q\in C_{c}^{1}(\Omega)^{d},|q(x)|_{2}\leq 1\text{ for allmost all }x\in\Omega\right\},$ (2)
@@
We choose to replace the solution space $BV(\Omega)\cap L^{2}(\Omega)$ in (1) by a set of neural networks $\mathcal{H}^{M}$, whose activation functions are ReLUs (rectifier linear units) and $M\in\mathbb{N}$ denotes the number of weights and biases of a neural network in $\mathcal{H}^{M}$. In the sequel we will call such neural networks ReLU neural networks (ReLU-NNs). While any other type of neural networks might be used in this approach too, we note that ReLU-NNs are theoretically quite well studied, which is useful for our theoretical analysis. The resulting optimization problem reads then
$\inf_{u\in\mathcal{H}^{M}}\alpha_{1}\|Tu-g\|_{L^{1}(\Omega)}+\alpha_{2}\|Tu-g\|_{L^{2}(\Omega)}^{2}+\lambda\operatorname{TV}(u).$ (3)
@@
As the Deep Ritz Method and PINNs, this approach can be interpreted as unsupervised learning. However, in a more classical view (3) is simply an optimization problem with solution space $\mathcal{H}^{M}$. Unfortunately, it turns out that even if (1) has a solution in $BV(\Omega)\cap L^{2}(\Omega)$ this solution is in general not in $\mathcal{H}^{M}$, see example 3.3 below. This has also been observed in [9]. The reason for this negative result is that the space $BV(\Omega)\cap L^{2}(\Omega)$ contains functions with jumps and ReLU-NNs cannot represent such functions. In fact ReLU-NNs can “only” represent continuous and piecewise affine functions [7]. Despite this negative observation, we suggest an auxiliary neural network model, which is a constrained version of (3). That is we add a constraint to (3) such that a solution in $\mathcal{H}^{M}$ is indeed guaranteed. We prove that this auxiliary model $\Gamma$-converges to the $L^{1}$-$L^{2}$-TV model, which fully justifies our approach. We are aware that $\Gamma$-convergence for a Deep Ritz Method is also studied in [21], but in a completely different setting, i.e. for a different type of optimization problem, namely for variational problems arising from partial differential equations, and in Sobolev spaces and not in the space of bounded variation. Hence our analysis completely differs from the one in [21]. However, our $\Gamma$-convergence proof is inspired by the $\Gamma$-convergence analysis in [34], where the $\Gamma$-convergence of a discrete total variation to the continuous total variation (2) is presented.
@@
To numerically compute a solution of the auxiliary model, numerical quadrature to approximate the integrals are needed. Note that since $\mathcal{H}^{M}\subset W^{1,1}(\Omega)$, we have that $\operatorname{TV}(u)=\int_{\Omega}|\nabla u|_{2}\text{d}x$ for any $u\in\mathcal{H}^{M}$. Therefore, the total variation of $u\in\mathcal{H}^{M}$ is computed using the gradient of $u$, which must be carefully approximated in a numerical setting, i.e. when applying quadrature. It seems reasonable to approximate the gradient by a finite difference approximation. The reason for this is that replacing the integral by a numerical quadrature evaluates the gradient only at a finite number of points. However, this can be problematic for piecewise constant functions, because the gradient may be zero at the specific evaluation points chosen by the quadrature, even though the overall function might have significant variation. This leads to an inaccurate computation of the total variation. A finite difference approximation of the gradient circumvents this shortcoming. We would like to emphasize that despite the described discretizations we still optimize the resulting discrete auxiliary model over the space $\mathcal{H}^{M}$. Further, we are able to show that this discrete auxiliary model again $\Gamma$-converges to (1). In particular, the $\Gamma$-convergence proof suggests a particular discretization of the gradient and hence total variation which ensures that the discrete solutions approximate the continuous solution as the discretization becomes finer. The same particular discretization of the total variation has been already used in [34] to show $\Gamma$-convergence and later in [33], but not in the context of neural networks. Our proof is actually based on the $\Gamma$-convergence proof in [34] and adjusted to our neural network setting. However, any convergence result of minimizers of the discrete auxiliary model to a solution of the continuous problem is missing in [34], while we present such a result in our setting. In [33] a convergence result of minimizers is presented, however with respect to the $L^{2}(\Omega)$ topology and under certain regularity assumption on the data $g$, which we do not need. Hence our result is more general and even stated with respect to the $L^{1}(\Omega)$ topology.
@@
The rest of the paper is organized as follows: In section 2, we review and introduce essential definitions related to neural networks and functional analysis, which are foundational for the subsequent sections. We analyze the neural network problem (3) in section 3 and in particular suggest an auxiliary model which indeed possesses a minimizer in $\mathcal{H}^{M}$. We show its $\Gamma$-convergence to the $L^{1}$-$L^{2}$-TV model and provide an error estimate. Moreover, we present a universal approximation result for $BV$-functions. In section 4 we describe the numerical realization of the proposed model. In particular, we present a numerical realization of the total variation. We connect the fully discretized neural network model to the original model (1) by utilizing $\Gamma$-convergence. Further we discuss a numerical issue and propose a solution to circumvent it. In section 5 we present several numerical experiments supporting our theoretical findings. We conclude with a short discussion in section 6.
@@ 2 Definitions and notations
As we are interested to represent functions $u:\Omega\to\mathbb{R}$ with a ReLU-NN, we consider only ReLU-NNs whose input layer consists of $d\in\{1,2\}$ neurons and whose output layer consists of 1 neuron. Thereby the number of neurons in the hidden layers might be arbitrary but fixed such that the number of weights and biases is $M\in\mathbb{N}$. We denote the set of ReLU-NNs with this fixed architecture by $\mathcal{H}^{M}$ to emphasize the number $M$ of weights and biases, which is in particularly descriptive in a convergence process where the number $M$ might change. The collection of all weights and biases of a neural network in $\mathcal{H}^{M}$ is denoted by $\theta\in\mathbb{R}^{M}$. For a neural network $u_{\theta}\in\mathcal{H}^{M}$ we use the index $\theta$ to emphasize that the neural network depends on the weights and biases $\theta$.
@@ 3 Problem description
$\min_{u_{\theta}\in\mathcal{H}^{M}}\alpha_{1}\|Tu_{\theta}-g\|_{L^{1}(\Omega)}+\alpha_{2}\|Tu_{\theta}-g\|_{L^{2}(\Omega)}^{2}+\lambda\operatorname{TV}(u_{\theta}).$ (4)
@@ 3.1 Existence of a minimizer
Let $\mathcal{H}^{M}$ be the set of one hidden layer neural networks, where each one hidden layer neural network is defined as
@@
Here $m\in\mathbb{N}$ is the number of neurons in the hidden layer, $w\in\mathbb{R}^{m}$ is the vector of weights from the input neuron to the hidden layer, $a\in\mathbb{R}^{m}$ is the vector of weights from the hidden layer to the output neuron, $b\in\mathbb{R}^{m}$ is a vector of biases, and $\sigma:\mathbb{R}\to\mathbb{R}$ is the ReLU activation function, i.e. $\sigma(x)=\max\{0,x\}$. Hence, we have that $M=3m$. If we try to approximate the BV-function
@@
on $\Omega=(-l_{\ell},l_{u})\subset\mathbb{R}$ by solving $\min_{u_{w,a,b}\in\mathcal{H}^{M}}\|u_{w,a,b}-g\|_{L^{2}(\Omega)}^{2}$, i.e. we set $T=I$, $\alpha_{1}=0=\lambda$ and $\alpha_{2}=1$ in (4), then it turns out that this problem does not have a solution. In fact, any function in $\mathcal{H}^{M}$ is continuous and since $g$ is not continuous we obtain that $\|u_{w,a,b}-g\|_{L^{2}(\Omega)}^{2}>0$ for all $u_{w,a,b}\in\mathcal{H}^{M}$. To see that this is not a contradiction to the universal approximation theorem, see theorems 3.1 and 3.2, i.e. that any $BV$-function can be approximated by a ReLU-NN arbitrarily good, let us set $m=2$, $a_{1}=1$ and $a_{2}=-1$. Further we choose $w_{1}=\frac{1}{h}=w_{2}$, $h>0$, and $b_{1}=1$, $b_{2}=0$. This yields
@@
where $c_{1},c_{2}\in\mathbb{R}^{+}_{0}$ such that $c_{1}\leq c_{2}$ and $c_{1},c_{2}$ depend on $\alpha_{1}$ and $\alpha_{2}$. The analytic computation of the exact solution for this example can be founded in appendix A. Hence, if $\alpha_{1}$ and $\alpha_{2}$ are chosen such that $c_{1}<c_{2}$, then for any $M\in\mathbb{N}$ a function in $\mathcal{H}^{M}$ cannot solve this problem, but only approximates a solution arbitrary close for sufficiently large $M$. From example 3.3 we learn that the failure of the existence of a solution of (4) might be due to the unboundedness of the weights $\theta$, rendering the solution space not compact. Thus we slightly modify (4) to
$\min_{u_{\theta}\in\mathcal{H}_{c}^{M}}\alpha_{1}\|Tu_{\theta}-g\|_{L^{1}(\Omega)}+\alpha_{2}\|Tu_{\theta}-g\|_{L^{2}(\Omega)}^{2}+\lambda\operatorname{TV}(u_{\theta})$ (9)
where $\mathcal{H}_{c}^{M}:=\{u_{\theta}\in\mathcal{H}^{M}\colon|\theta|_{\infty}\leq c\}$ with $c\geq 0$ typically to be chosen large such that (9) is a close approximation of (4). Note that in $\mathbb{R}^{M}$ all norms are equivalent and hence any arbitrary vector-norm could be chosen instead of $|\cdot|_{\infty}$.
@@
The optimization problem in (9) possesses a solution $u_{\theta}\in\mathcal{H}^{M}$.
@@
where $\Theta:=\{\theta\in\mathbb{R}^{M}\colon|\theta|_{\infty}\leq c\}$. Since $\Theta$ is closed and bounded, it is compact (Heine-Borel Theorem). Let us denote the functional in (10) by $J(\theta):=J_{c}(\theta)+J_{d}(\theta)$, where $J_{c}(\theta):=\alpha_{1}\|Tu_{\theta}-g\|_{L^{1}(\Omega)}+\alpha_{2}\|Tu_{\theta}-g\|_{L^{2}(\Omega)}^{2}$ and $J_{d}(\theta):=\lambda\operatorname{TV}(u_{\theta})$. We show now that $J$ is lower semi-continuous. Since the ReLU activation function $\sigma$ is continuous, functions in $\mathcal{H}^{M}$ are continuous with respect to $\theta$. Hence, for a sequence $(\theta_{n})_{n}\subset\mathbb{R}^{M}$ with limit $\theta\in\mathbb{R}^{M}$ for $n\to\infty$ we have $u_{\theta_{n}}\to u_{\theta}$ as $\theta_{n}\to\theta$. Since strong convergence implies weak convergence we also have $u_{\theta_{n}}\rightharpoonup u_{\theta}$ as $\theta_{n}\to\theta$. By the weak lower semi-continuity of the total variation we obtain $\operatorname{TV}(u_{\theta})\leq\liminf_{\theta_{n}\to\theta}\operatorname{TV}(u_{\theta_{n}})$ and hence $J_{d}$ is lower semi-continuous. Further, since $J_{c}$ is continuous, we have that $J$ is lower semi-continuous.
@@
We note that theorem 3.4 can be extended to any activation function $\sigma\in C(\mathbb{R})$, as the proof only requires that the activation functions are continuous. That is if in (9) the solution space $\mathcal{H}^{M}$ is replaced by a set of neural networks with arbitrary continuous activation functions, then (9) still possesses a solution in this respective space.
@@
$\min_{u_{\theta}\in\mathcal{H}^{M}}\alpha_{1}\|Tu_{\theta}-g\|_{L^{1}(\Omega)}+\alpha_{2}\|Tu_{\theta}-g\|_{L^{2}(\Omega)}^{2}+\lambda\operatorname{TV}(u_{\theta})+\alpha_{\theta}|\theta|_{\infty},$ (11)
@@
Note that theorem 3.4 guarantees “only” the existence of a minimizer, but does not say anything about the uniqueness. In fact, it seems elusive to show uniqueness of a solution of (9), since $u_{\theta}$ is usually non-convex with respect to $\theta$ rendering (9) non-convex. On the contrary (1) is convex with respect to $u$ and even strictly convex if $\alpha_{2}>0$ and $T$ is injective. This means that changing the solution space in (1) to $\mathcal{H}^{M}$ or $\mathcal{H}_{c}^{M}$ also changes the nature of the problem. Still (1) and (9) are closely related as we will show in the next subsection.